# Gradient Boosting Decision Trees on Medical Diagnosis over Tabular Data

### v2

@@ 1 Introduction
Motivated by these advancements, in this work, we highlight the superiority of the ensemble models, especially the GBDTs, over other traditional ML methods and DL architectures, in medical diagnosis tasks, specifically over tabular data. We compared 5 traditional ML methods and 5 DL models with 4 ensemble models, 3 of them being GBDT models, over 7 different medical datasets. We experimentally showed that GBDTs outperform all of the other methods and have the highest ranks over the medical diagnosis datasets on average. Unlike previous works that provide individual applications focusing on a limited number of disease types [7] or certain types of tabular datasets outside the medical field [2] by utilizing certain types of ML algorithms [8] or a limited number of comparisons regarding DL and ML methodologies [9], our contribution is to provide deep insights into the superiority of the GBDT models over state-of-the-art tabular DL and traditional ML models, focused on a diverse set of tabular medical diagnosis tasks consisting of different dataset dimensions and size settings. Having evaluated the relative performance of the methods across different clinical settings, we also provide insight into which model is more suitable in such medical cases. Additionally, we provide an analysis of training time consumption between the models, which has crucial implications in the medical field in terms of providing fast and accurate well-informed clinical decisions.
2 Background
2.1 Traditional Machine Learning Models
@@
2.2 Deep Neural Networks
@@
The development of DL algorithms is driven by the availability of medical datasets, which are often challenging to access due to their size and diversity. In other words, the proper acquisition of these datasets is integral for interpreting model performance and accuracy within real-world clinical data. Furthermore, LLM processing employs ethical concerns together with security and privacy due to the inclusion of personal health records during model training [30]. These concerns and challenges are broadly justified in the literature in terms of its acceptability [31], downstream accuracy [32], medical ethics [33], [34], [35]. Hence, due to the limitations mentioned above, few experimental studies of DL applications in medicine have been conducted, thus, there is a great demand for explanatory research to validate the the diverse models using medical data. On a larger scale, employing diverse deep neural network architectures can serve as a guiding tool for handling more complex data.
@@ 2.3 Ensemble Models
More recently, gradient-boosting methods have started to become the state-of-the-art in many fields, particularly GBDTs. Unlike other ensemble learning methods, many weak learners are combined to create a single strong learner, with each weak learner representing an individual decision tree. Each tree tries to minimize the error on the previous tree, where all of the trees are connected sequentially. Therefore, the model improves at each step iteratively, resulting in a stronger learner by the end. Some of their application fields include GPS signal reception classification [45] and credit risk assessments [46], and also occupy various applications in the medical field including heart disease detection [47], and parkinson’s disease progression prediction [48]. GBDT comprises three main algorithms that have been occupying the field in recent years, which are XGBoost [49], LightGBM [50], and CatBoost [51].
@@ 3 Problem Formulation
Several techniques can be used in order to overcome possible challenges, including imbalanced datasets and generalization to unseen data, that will be mentioned in Section 4.
@@
Table 1: Dataset information. The number of features includes the target labels as well.
Datasets Source Samples Features Classes Task
CD [53] OpenML 70k 12 2 Binary
Heart Failure [54] Kaggle 299 13 2 Binary
Parkinsons [55] Kaggle 195 23 2 Binary
EEG Eye State [56] OpenML 15k 15 2 Binary
Eye Movements [57] OpenML 11k 28 3 Multi-Class
Arcene [58] OpenML 200 10k 2 Binary
Prostate [59] OpenML 102 12.6k 2 Binary
4 Experiments
4.1 Datasets
@@
We evaluated the performance of GBDT models over 7 medical diagnosis datasets. Dataset information that includes the number of samples, features, classes, and the type of the task is provided in Table 1. Cardiovascular Disease Dataset (CD) [53] contains heart disease data that consist of 70000 patient records with 11 independent features, including a binary label. The aim is to build a predictive algorithm that detects the presence of an early-stage heart disease. Heart Failure [54] is another heart-related dataset that contains 299 medical records of different patients who had heart failure, where each of the patients has 13 clinical features. The aim is to predict the death event for the given patient records, i.e., whether there will occur heart failure with respect to the given patient features. Parkinson’s [55] dataset comprises several biomedical voice measurements from a total of 31 patients, where 23 of them have Parkinson’s disease. There is a total of 195 different samples from these patients, where each of them includes 22 features. These features include continuous measurements. It is used to predict the status of a given patient data, which corresponds to being healthy or having Parkinson’s disease. EEG Eye State [56] dataset consists of 14 features for each patient that hold values regarding the EEG measurements. These features include continuous measurements. There are nearly 15000 samples and each of them has a label indicating the eye-closed or eye-open states. It is aimed to predict the state of a patient sample using these EEG measurement features. Eye Movements [57] contains 27 features in which 22 of them comprises features that are commonly used in psychological studies on eye movements. The rest of the features include extra information about the samples, that each of the samples is an assignment. These assignments consist of a question, that is followed up by several answer sentences. The task is to predict the multi-class relevance label of an answer for each assignment, in which the labels are irrelevant, relevant, and the correct answer. Arcene [58] dataset includes several features that indicate the abundance of proteins in human sera, having a given mass value. There are 200 patient samples and nearly 10000 features in total. These features are continuous. The task is to predict patients who have cancer and patients that are healthy, according to the given features. Lastly, Prostate [59] dataset includes continuous features that give information regarding prostate cancer. There are 102 patient samples and nearly 12600 total features. The aim is to predict whether a given patient has prostate cancer or not.
Table 2: Optimal Parameters for MLP
@@
Table 3: Optimal Parameters for STG
@@
Table 4: Optimal Parameters for TabNet
@@
Table 5: Optimal Parameters for TabTransformer
@@
Table 6: Optimal Parameters for VIME
@@ 4.2 Data Preprocessing
4.3 Training Procedure
We compared 5 traditional ML methods, which are SVM, Logistic Regression, KNN, Decision Tree, and Linear Discriminant Analysis (LDA), and 5 DL models, which are Multilayer Perceptron (MLP), STG, TabNet, TabTransformer, and VIME, with 4 ensemble models including Random Forest, 3 of them being GBDT models that are XGBoost, LightGBM, and CatBoost, over the different medical datasets in Table 1.
We selected ROC AUC score as our comparison metric. ROC AUC score is the area under the ROC curve that summarizes the the classifier performance with different decision thresholds, and it is a plot of true positive rate (TPR) against the false positive rate (FPR). It is a commonly used metric in comparing multiple models due to its drift tolerance in class balances and reliability in imbalanced data. Hence, we utilized ROC AUC score for both the optimization of the hyperparameters, which is explained below at section 4.4, and the comparison of different models for each dataset.
@@ 4.3 Training Procedure
4.4 Hyperparameter Optimization
Table 7: ROC AUC scores for the traditional ML models, state-of-the-art tabular DL methods, and GBDT models on different datasets. Higher is better. The best is bolded, and the second best is underlined.
@@ 4.4 Hyperparameter Optimization
Table 8: Different metric scores for highest ranked (LightGBM) model on different datasets.
@@
Traditional ML algorithms and ensemble models are easier to optimize compared to deep neural networks due to their less complex architectures. For the deep learning models, each hyperparameter combination setup is trained for 1000 epochs with an early stopping criterion with 100 epochs patience. On average, 36 different combinations are evaluated for each model, in which the best combination is taken as the best performance of the specific model for comparison. Optimized hyperparameters for the DL architectures with respect to the datasets are shown in Table’s 2 - 6.
4.5 Results
Overall results are shown in Table 7. Best models are selected by comparing the average ROC AUC scores of the 8-folds over each respective dataset. For the multi-class problem in the Eye Movements dataset, ’one-vs-one’ configuration of the ROC AUC scores are reported. Each result also includes the standard deviation between the different fold results. According to the performances, each model is ranked with respect to the results on each of the datasets, and their average ranks are reported at the end of the table. It can be inferred that GBDT consistently outperforms both the traditional ML methods and the state-of-the-art tabular DL models regardless of the dataset sample size. They perform well in both small sized (e.g. Prostate) and large sized (e.g. CD) datasets. Top-3 ranks consist of the three proposed GBDT models, which demonstrate the overall superior performance of the GBDT over the other traditional ML and tabular DL methods in medical diagnosis tasks.
@@ 4.5 Results
Consequently, different metric scores for the highest ranked GBDT model in average, which is LightGBM, are shown in Table 8. Metrics are Accuracy, F1 Score, ROC AUC Score (same as in Table 7), Precision, and Recall. Again, the average and standard deviation of the fold scores are reported. Respective formulas for the metrics are represented at Eq’s (3) - (6) except ROC AUC, which is defined in Section 4.4. TP, FP, TN, and FN are defined as true positives, false positives, true negatives, and false negatives respectively. Regarding the multi-class problem in the Eye Movements dataset, weighted averages of the label scores are reported for F1 Score, Recall, and Precision metrics.
@@
(a) Eye Movements dataset
(b) Prostate dataset
5 Discussion
Observing Table 7, traditional ML models such as SVM, Logistic Regression, and KNN’s generally show varying performance on different tasks, such as KNN excelling on Parkinson’s but struggling on others. The reason behind is that SVM and logistic regression are linear models that usually rely on low-dimensional datasets to perform well. Since medical diagnosis datasets usually consist of diverse patient symptoms, such linear models may struggle in performance. On the other hand, being a distance-based method, KNN can perform inconsistently and be sensitive to noisy and high-dimensioned data, which is usually the case for datasets in the medical field. Unlike other ML models, Random Forest provides accurate performances over both small and large sample-sized datasets, such as EEG Eye State and Heart Failure, since it is an ensemble method that becomes effective in reducing overfitting and capturing non-linear relations, being a preferable algorithm for medical datasets with complex patterns. Furthermore, GBDT models, in which we evaluated LightGBM, CatBoost, and XGBoost, tend to perform better on tabular data. These models effectively capture complex interactions between the features that become advantageous for handling imbalanced datasets. In medical settings where the data includes noisy features, GBDT possesses robust mechanisms to overcome these difficulties more easily than other ML methodologies.

### v3

@@ I Introduction
Motivated by these advancements, in this work, we highlight the superiority of the ensemble models, especially the GBDTs, over other traditional ML methods and DL architectures, in medical diagnosis tasks, specifically over tabular data. We compared 5 traditional ML methods and 5 DL models with 4 ensemble models, 3 of them being GBDT models, over 7 different medical datasets. We experimentally showed that GBDTs outperform all of the other methods and have the highest ranks over the medical diagnosis datasets on average. Unlike previous works that provide individual applications focusing on a limited number of disease types [7] or certain types of tabular datasets outside the medical field [2] by utilizing certain types of ML algorithms [8] or a limited number of comparisons regarding DL and ML methodologies [9], our contribution is to provide deep insights into the superiority of the GBDT models over state-of-the-art tabular DL and traditional ML models, focused on a diverse set of tabular medical diagnosis tasks consisting of different dataset dimensions and size settings. Having evaluated the relative performance of the methods across different clinical settings, we also provide insight into which model is more suitable in such medical cases. Additionally, we provide an analysis of training time consumption between the models, which has crucial implications in the medical field in terms of providing fast and accurate well-informed clinical decisions. Our code is publicly available [^1] https://github.com/yarkin06/TabularGBDT.
II Background
II-A Traditional Machine Learning Models
@@
II-B Deep Neural Networks
@@
The development of DL algorithms is driven by the availability of medical datasets, which are often challenging to access due to their size and diversity. In other words, the proper acquisition of these datasets is integral for interpreting model performance and accuracy within real-world clinical data. Furthermore, LLM processing employs ethical concerns together with security and privacy due to the inclusion of personal health records during model training [30]. These concerns and challenges are broadly justified in the literature in terms of its acceptability [31], downstream accuracy [32], medical ethics [33], [34]. Hence, due to the limitations mentioned above, few experimental studies of DL applications in medicine have been conducted, thus, there is a great demand for explanatory research to validate the the diverse models using medical data. On a larger scale, employing diverse deep neural network architectures can serve as a guiding tool for handling more complex data.
@@ II-C Ensemble Models
More recently, gradient-boosting methods have started to become the state-of-the-art in many fields, particularly GBDTs. Unlike other ensemble learning methods, many weak learners are combined to create a single strong learner, with each weak learner representing an individual decision tree. Each tree tries to minimize the error on the previous tree, where all of the trees are connected sequentially. Therefore, the model improves at each step iteratively, resulting in a stronger learner by the end. Some of their application fields include GPS signal reception classification [44], radar waveform classification [45], and credit risk assessments [46]. Moreover, they occupy various applications in the medical field including heart disease detection [47] and parkinson’s disease progression prediction [48]. GBDT comprises three main algorithms that have been occupying the field in recent years, which are XGBoost [49], LightGBM [50], and CatBoost [51].
@@ III Problem Formulation
Several techniques can be used in order to overcome possible challenges, including imbalanced datasets and generalization to unseen data, that will be mentioned in Section IV.
@@
TABLE I: Dataset information. The number of features includes the target labels as well.
Datasets Samples Features Classes Task
CD [53] 70k 12 2 Binary
Heart Failure [54] 299 13 2 Binary
Parkinsons [55] 195 23 2 Binary
EEG Eye State [56] 15k 15 2 Binary
Eye Movements [57] 11k 28 3 Multi-Class
Arcene [58] 200 10k 2 Binary
Prostate [59] 102 12.6k 2 Binary
IV Experiments
IV-A Datasets
@@
We evaluated the performance of GBDT models over 7 medical diagnosis datasets. Dataset information that includes the number of samples, features, classes, and the type of the task is provided in Table I. Cardiovascular Disease Dataset (CD) [53] contains heart disease data that consist of 70000 patient records with 11 independent features, including a binary label. The aim is to build a predictive algorithm that detects the presence of an early-stage heart disease. Heart Failure [54] is another heart-related dataset that contains 299 medical records of different patients who had heart failure, where each of the patients has 13 clinical features. The aim is to predict the death event for the given patient records, i.e., whether there will occur heart failure with respect to the given patient features. Parkinson’s [55] dataset comprises several biomedical voice measurements from a total of 31 patients, where 23 of them have Parkinson’s disease. There is a total of 195 different samples from these patients, where each of them includes 22 features. These features include continuous measurements. It is used to predict the status of a given patient data, which corresponds to being healthy or having Parkinson’s disease. EEG Eye State [56] dataset consists of 14 features for each patient that hold values regarding the EEG measurements. These features include continuous measurements. There are nearly 15000 samples and each of them has a label indicating the eye-closed or eye-open states. It is aimed to predict the state of a patient sample using these EEG measurement features. Eye Movements [57] contains 27 features in which 22 of them comprises features that are commonly used in psychological studies on eye movements. The rest of the features include extra information about the samples, that each of the samples is an assignment. These assignments consist of a question, that is followed up by several answer sentences. The task is to predict the multi-class relevance label of an answer for each assignment, in which the labels are irrelevant, relevant, and the correct answer. Arcene [58] dataset includes several features that indicate the abundance of proteins in human sera, having a given mass value. There are 200 patient samples and nearly 10000 features in total. These features are continuous. The task is to predict patients who have cancer and patients that are healthy, according to the given features. Lastly, Prostate [59] dataset includes continuous features that give information regarding prostate cancer. There are 102 patient samples and nearly 12600 total features. The aim is to predict whether a given patient has prostate cancer or not.
TABLE II: Optimal Parameters for MLP
@@
TABLE III: Optimal Parameters for STG
@@
TABLE IV: Optimal Parameters for TabNet
@@
TABLE V: Optimal Parameters for TabTransformer
@@
TABLE VI: Optimal Parameters for VIME
@@ IV-B Data Preprocessing
IV-C Training Procedure
We compared 5 traditional ML methods, which are SVM, Logistic Regression, KNN, Decision Tree, and Linear Discriminant Analysis (LDA), and 5 DL models, which are Multilayer Perceptron (MLP), STG, TabNet, TabTransformer, and VIME, with 4 ensemble models including Random Forest, 3 of them being GBDT models that are XGBoost, LightGBM, and CatBoost, over the different medical datasets in Table I.
We selected ROC AUC score as our comparison metric. ROC AUC score is the area under the ROC curve that summarizes the the classifier performance with different decision thresholds, and it is a plot of true positive rate (TPR) against the false positive rate (FPR). It is a commonly used metric in comparing multiple models due to its drift tolerance in class balances and reliability in imbalanced data. Hence, we utilized ROC AUC score for both the optimization of the hyperparameters, which is explained below at section IV-D, and the comparison of different models for each dataset.
@@ IV-C Training Procedure
IV-D Hyperparameter Optimization
TABLE VII: ROC AUC scores for the traditional ML models, state-of-the-art tabular DL methods, and GBDT models on different datasets. Higher is better. The best scores are bolded, and the second bests are underlined.
@@ IV-D Hyperparameter Optimization
TABLE VIII: Different metric scores for highest ranked (LightGBM) model on different datasets.
@@
Traditional ML algorithms and ensemble models are easier to optimize compared to deep neural networks due to their less complex architectures. For the deep learning models, each hyperparameter combination setup is trained for 1000 epochs with an early stopping criterion with 100 epochs patience. On average, 36 different combinations are evaluated for each model, in which the best combination is taken as the best performance of the specific model for comparison. Optimized hyperparameters for the DL architectures with respect to the datasets are shown in Table’s II - VI.
IV-E Results
Overall results are shown in Table VII. Best models are selected by comparing the average ROC AUC scores of the 8-folds over each respective dataset. For the multi-class problem in the Eye Movements dataset, ’one-vs-one’ configuration of the ROC AUC scores are reported. Each result also includes the standard deviation between the different fold results. According to the performances, each model is ranked with respect to the results on each of the datasets, and their average ranks are reported at the end of the table. It can be inferred that GBDT consistently outperforms both the traditional ML methods and the state-of-the-art tabular DL models regardless of the dataset sample size. They perform well in both small sized (e.g. Prostate) and large sized (e.g. CD) datasets. Top-3 ranks consist of the three proposed GBDT models, which demonstrate the overall superior performance of the GBDT over the other traditional ML and tabular DL methods in medical diagnosis tasks.
@@ IV-E Results
Consequently, different metric scores for the highest ranked GBDT model in average, which is LightGBM, are shown in Table VIII. Metrics are Accuracy, F1 Score, ROC AUC Score (same as in Table VII), Precision, and Recall. Again, the average and standard deviation of the fold scores are reported. Respective formulas for the metrics are represented at Eq’s (3) - (6) except ROC AUC, which is defined in Section IV-D. TP, FP, TN, and FN are defined as true positives, false positives, true negatives, and false negatives respectively. Regarding the multi-class problem in the Eye Movements dataset, weighted averages of the label scores are reported for F1 Score, Recall, and Precision metrics.
@@
V Discussion
Observing Table VII, traditional ML models such as SVM, Logistic Regression, and KNN’s generally show varying performance on different tasks, such as KNN excelling on Parkinson’s but struggling on others. The reason behind is that SVM and logistic regression are linear models that usually rely on low-dimensional datasets to perform well. Since medical diagnosis datasets usually consist of diverse patient symptoms, such linear models may struggle in performance. On the other hand, being a distance-based method, KNN can perform inconsistently and be sensitive to noisy and high-dimensioned data, which is usually the case for datasets in the medical field. Unlike other ML models, Random Forest provides accurate performances over both small and large sample-sized datasets, such as EEG Eye State and Heart Failure, since it is an ensemble method that becomes effective in reducing overfitting and capturing non-linear relations, being a preferable algorithm for medical datasets with complex patterns. Furthermore, GBDT models, in which we evaluated LightGBM, CatBoost, and XGBoost, tend to perform better on tabular data. These models effectively capture complex interactions between the features that become advantageous for handling imbalanced datasets. In medical settings where the data includes noisy features, GBDT possesses robust mechanisms to overcome these difficulties more easily than other ML methodologies.